OpenStudy (mitu12):

the perimeter of a rectangle is 52. the length is 6 greater that the width . find the length and width of this rectangle

OpenStudy (mitu12):

@tester97

OpenStudy (anonymous):

length=16 width=10

OpenStudy (tester97):

2L + 2W = 52 L + W = 26 L = W + 6W = 7W L + W = 26 7W + W = 26 8W = 26 W = 13/4 L = 91/4

OpenStudy (mitu12):

im confused

OpenStudy (anonymous):

a sec

OpenStudy (anonymous):